# Autoregressive Models in Vision: A Survey

### v1

@@ 3.3.3 Scene Generation
Inspired by the success of multi-scale alignment of visual and language tokens, more tasks are adopting similar approaches. Make-A-Scene (Gafni et al., 2022) combines text descriptions with scene layouts to generate images, improving control and alignment with human visual preferences through segmentation-based conditioning. SceneScript (Avetisyan et al., 2024) proposes a novel autoregressive image generation approach by using structured language commands to represent scenes. This method provides compact and interpretable scene representations, excelling in architectural layout estimation and 3D object detection, while being easily extendable to new tasks.
@@ 3.4.2 The Framework for Unifying Multimodal Understanding and Generation
The unified framework extends traditional multimodal understanding (Li et al., 2023a; Zhang et al., 2023b; Shi et al., 2023; Xie et al., 2024b; Wan et al., 2024) by enabling both visual and textual output generation. Unlike earlier models that focused solely on interpretation, this framework leverages the integration of Large Language Models (LLMs) to enhance multimodal content generation. The overview of auto-regressive multimodal LLM’s application is shown in Figure 10.
@@
With the rise of LLMs, autoregressive models have increasingly interfaced with diffusion models to enhance multi-modal capabilities. NEXT-GPT (Wu et al., 2023c) exemplifies this trend by combining the autoregressive approach with diffusion processes to improve image synthesis from textual inputs. The SEED (Ge et al., 2024) and EMU-series  (Sun et al., 2024b; 2023; Wang et al., 2024d) further explore this synergy, utilizing diffusion models to refine and stabilize the output quality of generated visual content. LaViT (Jin et al., ) employs a token merging strategy, which reduces the number of visual tokens by merging tokens that are spatially or semantically similar, thereby lowering computational costs while maintaining high-quality visual generation. Video-LaViT (Jin et al., 2024b), an extension of LaViT, applies a keyframe-based token reduction technique to video data, selecting keyframes that capture the most critical information and sparsely sampling intermediate frames, significantly reducing token count while preserving temporal coherence in video understanding and generation. Additionally, models like X-ViLA (Ye et al., 2024) (Cross-modality Vision, Language, Audio) push the boundaries of multi-modal learning by serving as a foundation model for cross-modality understanding, reasoning, and generation across video, image, language, and audio domains. X-ViLA demonstrates the potential of unified multi-modal architectures to handle diverse tasks across multiple data types, offering a holistic approach to multi-modal content generation. This integration has proven beneficial in addressing the limitations of purely autoregressive methods, particularly in generating high-fidelity and diverse images.
@@ 5.0.0.2 Discrete or continuous?
Autoregressive models have traditionally been associated with discrete representations. However, given that visual data are inherently continuous, most approaches require an additional discretization step, which, as discussed earlier, is far from straightforward. Recent studies (Li et al., 2024c) argue that the essence of autoregressive models lies in "next element prediction", regardless of whether the elements are discrete or continuous. This perspective revives the debate over the merits of “discrete vs. continuous” representations in autoregressive models. Continuous representations offer advantages, particularly in simplifying the training of visual data compressors. However, they also pose new challenges for designing autoregressive architectures. Firstly, the cross-entropy loss is no longer applicable. Although L2 loss is a simple alternative, it often compromises the quality and diversity of generated outputs (Li et al., 2024c). Thus, the development of alternative loss functions tailored for continuous settings is still under exploration. Secondly, the challenge of multimodal adaptability arises. Large language models (LLMs) based on discrete representations have already shown remarkable capabilities, but integrating visual continuous and discrete language representations within a unified autoregressive framework remains a complex task. TransFusion (Zhou et al., 2024) has made pioneering attempts in this area, yet significant work remains to build upon these initial efforts. Currently, continuous visual representations have yet to demonstrate a decisive advantage over discrete ones in autoregressive modeling. The full potential of autoregressive model-based continuous representations remains to be realized and represents a promising direction for future research.

### v2

@@ 3.3.3 Scene Generation
Inspired by the success of multi-scale alignment of visual and language tokens, more tasks are adopting similar approaches.SceneScript (Avetisyan et al., 2024) proposes a novel autoregressive image generation approach by using structured language commands to represent scenes. This method provides compact and interpretable scene representations, excelling in architectural layout estimation and 3D object detection, while being easily extendable to new tasks.
@@ 3.4.1 The Framework for Multimodal Understanding
Several works like the AIM series (El-Nouby et al., 2024; Fini et al., 2024) have also explored a compelling framework for learning visual representations through generative pre-training. In AIM (El-Nouby et al., 2024), images are decomposed into non-overlapping patches, and a causal Vision Transformer (ViT) (Dosovitskiy, 2020) is trained to autoregressively predict the next patch’s representation. These predicted representations are then mapped to pixel space via a heavyweight MLP for pixel-level reconstruction. This approach enables AIM to learn robust visual representations using only a generative objective. AIMV2 Fini et al. (2024) further enhances this by aligning visual representations with textual features, expanding its applicability to multimodal understanding tasks. Both models rely solely on generative pretraining and exhibit strong scaling properties, mirroring trends observed in LLMs Achiam et al. (2023); Dubey et al. (2024), and demonstrating that large autoregressive models can serve as a versatile paradigm for unified representation learning across modalities.
@@ 3.4.2 The Framework for Unifying Multimodal Understanding and Generation
The unified framework extends traditional multimodal understanding (Chen et al., 2022; Li et al., 2023a; Zhang et al., 2023b; Shi et al., 2023; Xie et al., 2024b; Wan et al., 2024) by enabling both visual and textual output generation. Unlike earlier models that focused solely on interpretation, this framework leverages the integration of Large Language Models (LLMs) to enhance multimodal content generation. The overview of auto-regressive multimodal LLM’s application is shown in Figure 10.
@@
With the rise of LLMs, autoregressive models have increasingly interfaced with diffusion models to enhance multi-modal capabilities. NEXT-GPT (Wu et al., 2023c) exemplifies this trend by combining the autoregressive approach with diffusion processes to improve image synthesis from textual inputs. The SEED (Ge et al., 2024) and EMU-series  (Sun et al., 2024b; 2023; Wang et al., 2024d) further explore this synergy, utilizing diffusion models to refine and stabilize the output quality of generated visual content. LaViT (Jin et al., ) employs a token merging strategy, which reduces the number of visual tokens by merging tokens that are spatially or semantically similar, thereby lowering computational costs while maintaining high-quality visual generation. Video-LaViT (Jin et al., 2024c), an extension of LaViT, applies a keyframe-based token reduction technique to video data, selecting keyframes that capture the most critical information and sparsely sampling intermediate frames, significantly reducing token count while preserving temporal coherence in video understanding and generation. Additionally, models like X-ViLA (Ye et al., 2024) (Cross-modality Vision, Language, Audio) push the boundaries of multi-modal learning by serving as a foundation model for cross-modality understanding, reasoning, and generation across video, image, language, and audio domains. X-ViLA demonstrates the potential of unified multi-modal architectures to handle diverse tasks across multiple data types, offering a holistic approach to multi-modal content generation. This integration has proven beneficial in addressing the limitations of purely autoregressive methods, particularly in generating high-fidelity and diverse images.
@@ 4.2 Evaluation of Visual Autoregressive Generation
4.3 Task-Specific Evaluation Metrics
In the previous sections, we categorized and introduced several common metrics based on various evaluation dimensions. In this subsection, we reorganize these metrics according to specific tasks. The aim is to provide task-specific classifications, allowing readers to quickly identify the most relevant metrics for each task. Specifically, metrics for visual autoregressive models are categorized as follows:
  • Reconstruction
    • PSNR, SSIM (Wang et al., 2004), LPIPS (Zhang et al., 2018), rFID (Heusel et al., 2017) are pixel-level accuracy metrics for restoration tasks, which is commonly adopted for evaluating the performance of tokenizers in token-based autoregressive models. Additionally, codebook utilization is often employed to assess how effectively vectors in VQ codebooks are utilized.
  • Image Generation
    • General Metrics. FID (Heusel et al., 2017), IS (Salimans et al., 2016), and KID (Bińkowski et al., 2018) are widely adopted for evaluating the quality of generated images. Precision and Recall (Kynkäänniemi et al., 2019) are typically used to assess the fidelity and diversity of generated content.
    • Text-to-Image Generation. CLIPSim (Radford et al., 2021) and R-Precision (Craswell, 2009) are key metrics for evaluating the relevance between text and images in text-to-image generation tasks.
    • Inpainting In addition to the general quality metrics, pixel-level metrics are employed to assess whether the inpainting results align with the ground truth.
  • Video Generation
    • General Metrics. FVD (Unterthiner et al., 2018) is commonly used to evaluate the video quality. CLIPSIM-Temp (Radford et al., 2021) and Wrapping Errors (Lai et al., 2018) are used to assess the temporal coherence across video generation tasks.
    • Text-to-Video Generation. Similar to T2I Generation, T2V generation uses semantic consistency metrics like CLIPSIM to evaluate alignment with given text descriptions. Benchmarks such as VBench (Huang et al., 2024) and EvalCrafter (Liu et al., 2024h) are also widely used to evaluate the overall performance of T2V generation.
  • Task-Specific Metrics
    • In addition, there are various task-specific metrics and benchmarks tailored to particular tasks. These are designed to assess specified attributes according to task requirements. For example, T2I-CompBench (Huang et al., 2023a) is used for evaluating compositional generation tasks, while CLIPSIM-Img is applied to example-based image editing tasks.
@@ 5.1.0.2 Discrete or continuous?
Autoregressive models have traditionally been associated with discrete representations. However, given that visual data are inherently continuous, most approaches require an additional discretization step, which, as discussed earlier, is far from straightforward. Recent studies (Li et al., 2024c) argue that the essence of autoregressive models lies in "next element prediction", regardless of whether the elements are discrete or continuous. This perspective revives the debate over the merits of “discrete vs. continuous” representations in autoregressive models. Continuous representations offer advantages, particularly in simplifying the training of visual data compressors. However, they also pose new challenges for designing autoregressive architectures. Firstly, the cross-entropy loss is no longer applicable. Although L2 loss is a simple alternative, it often compromises the quality and diversity of generated outputs (Li et al., 2024c). Thus, developing alternative loss functions tailored for continuous settings is still under exploration. Secondly, the challenge of multimodal adaptability arises. Large language models (LLMs) based on discrete representations have already shown remarkable capabilities, but integrating visual continuous and discrete language representations within a unified autoregressive framework remains a complex task. TransFusion (Zhou et al., 2024) has made pioneering efforts in this area, yet significant work remains to build upon these initial efforts. Currently, continuous visual representations have yet to demonstrate a decisive advantage over discrete ones in autoregressive modeling. The full potential of autoregressive model-based continuous representations remains to be realized and represents a promising direction for future research.
@@
5.2 Application Roadmaps
5.2.0.1 Long Video Generation
The generation of long videos remains an unsolved problem for various generative models. Due to the inherent limitations of current generation algorithms, the most advanced flow-based video generation models (OpenAI, 2024; Kuaishou, 2024) typically produce videos with durations similar to the training data, usually around 5 to 10 seconds. However, autoregressive models, particularly in the context of large language models (LLMs), have demonstrated strong extrapolation abilities, generating text sequences that far exceed the length of the sequences seen during training. This capability suggests that visual autoregressive models hold significant potential for generating long video sequences. Recent work (Jin et al., 2024b) has begun to explore the use of autoregressive models for generating longer video sequences, opening up new avenues for this challenging task.
5.2.0.2 World Simulator for Embodied AI
A promising direction for the future of visual autoregressive generation is the development of world simulators that can facilitate the training of models in dynamic, interactive environments. These simulators would enable models to generate realistic visual content by interacting with simulated worlds in a context-aware manner. By integrating physical dynamics, environmental factors, and agent interactions, such simulators could allow visual autoregressive models to generate long and temporally consistent video sequences, building on the model’s general ability to extrapolate from sufficient training data. Existing works (Hu et al., 2023; Bruce et al., 2024; Wu et al., 2024b; Cheang et al., 2024) have already made significant progress in leveraging autoregressive models as world models. The potential to scale training data and model size to enable the learning of general intelligence, allowing models to interact with the world and provide multimodal feedback, presents an exciting avenue for further exploration.
5.2.0.3 Unified Multimodal Generation.
The future of multimodal autoregressive generation lies in advancing models capable of seamlessly combining and generating content across different modalities, such as text, images, audio, and video. Autoregressive models naturally lend themselves to cross-modal generation, as they treat different modalities equally in the token space. While existing work has explored the potential of autoregressive models for multimodal generation (Lu et al., 2022; Wang et al., 2024d) and unified models for both understanding and visual generation (Team, 2024; Zhou et al., 2024; Xie et al., 2024b; Wu et al., 2024a), these models still fall short of the state-of-the-art models in each individual modality. The key challenge likely lies in the development of large-scale multimodal datasets and unified cross-modal representations. We hope an aha moment that, with sufficient training, autoregressive models will emerge with powerful cross-modal generation capabilities, where the strengths of one modality can boost the generation of others.
@@ 6 Conclusion
7 Acknowledgement
This work was primarily supported by the Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC), Hong Kong SAR, and partially supported by ONR, NSF, the Simons Foundation, as well as gifts and awards from Google, Amazon, and Apple.